# Diagnostic Reasoning in Natural Language: Computational Model and Application

### v1

@@ A.1.1 Interviews and Formative Study
First, there are two general types of papers in biomedicine relevant to their work: natural science papers, which investigate natural phenomena, and method papers, which propose a new approach to conducting experiments. Second, biomedical papers follow a relatively rigid surface structure with an introduction including background information, the methods section as a detailed experimental log, the results section reporting the exact measurements, and the discussion section followed by an overall conclusion. The results and discussion section can be inter-mingled. Third, figures and tables play a major role in the understanding and assessment of a paper. Fourth, methodological soundness and consistency of results are key indicators during text assessment. Finally, several heuristics exist to detect bad experiment design, such as the lack of controls or suspiciously high standard deviations on measurements.
@@
We conducted the interviews with each expert in isolation and orienting towards best practices of interview design by Knott et al. (2022). To structure the interview, we discussed the assessment criteria and process based on reviewing checklists of the field, starting with a general and intuitive description of the assessment process and increasing the level of detail. For the first interview, we considered the manuscript checklist by Seals and Tanaka (2000) conceptualized for physiology research. For the second interview, we use the Portland Press review checklist [^6] https://portlandpress.com/pages/peer_review_checklist (accessed 08/2024) for biomedical sciences. We opted for two different checklists to encourage different perspectives during the workflow elicitation and mitigate the impact of the interviewer’s biases.
The result of these interviews is two different step-wise manuals on how to iterate through a biomedical paper and assess its methodological soundness. We identify the there types of actions taken during assessment (read, extract, infer) which align between both manuals. While some of the subordinate steps deviate and the researchers put a different emphasize on each of these aspects, the overall assessment approach is very similar. In summary, first, the whole article is skimmed to get an understanding of its contents. Then, the experts check the contextualization with related work and background information in the introduction. Afterward, they skip to the results section, which includes figures and tables. They check the alignment of results and whether the reporting seems plausible and follows best scientific practice. Subsequently, they read the discussion in detail and analyze if the provided interpretations and conclusions align with their own readings of the results and their context in related work. Then, they verify that the method section covers all important methods and that they are described correctly. Finally, they check the conclusion section, abstract and title to verify that they align with findings and methodology described in the other sections. Both experts agree on the temporal ordering of these major steps during assessment.
@@
After the interviews, we consolidate the sequential manuals into one manual covering all the mentioned points by both researchers. We then interpret the temporal ordering and the sub-ordination relationship between major and minor steps as causal dependencies to answer follow-up questions. For instance, to answer a sub-ordinate step, "Is the description of the state-of-the-art complete?" we require the input from the super-ordinate step, "How do the authors describe the state-of-the-art in the field?".
@@
First, the annotators did not finish the full workflow of $106$ steps within the three hours of the study, suggesting that the workflow should be subselected to keep the annotation practical. Second, the annotators tended to ignore the causal dependencies between the steps, meaning that they answered several steps in isolation based on the paper and not their own prior answers. The main obstacle for this lied in the lack of a tailored annotation interface that presents prior answers and enforces the causal structure. Third, annotators need training to get an overview of the workflow graph to avoid redundantly stating similar facts across multiple steps. An easy option to revise prior answers might also be interesting for that purpose.
@@
A.1.2 SCM for Biomedical Paper Assessment
@@
The full causal graph of the biomedical workflow is illustrated in Figure 5. Tables 4 to 6 list all workflow steps, including their type and prompt. The description and example are omitted for readability. We defer the reader to the associated codebase for the complete workflow step specifications.
@@
A.2 Data Collection
A.2.1 Underlying Paper Dataset
@@
A.2.2 Annotation Study Details
@@
A.3 Data Insights
A.3.1 Detailed Analysis of Answer Variability
@@
A.3.2 Structural Equation Approximation from Data
@@ A.4.1 Experimental Configurations
We design two prompts for each model, one for the extract steps and one for the infer steps. Table 10 shows the prompts for LLama-3, Table 11 for Mixtral, Table 8 for GPT-3.5-turbo, and Table 9 for GPT-4o. The template parameters are substituted before inference: the parent parameter is replaced by the paper text or prior answers in the graph; the task parameter is replaced by the prompt for that step in the workflow; the description is replaced by the detailed task description of that step in the workflow; the format specifies a json object matching the expected output type (e.g. including a boolean field for boolean questions); example is substituted by the same example that annotators saw during annotation but in a suitable json format.
@@
A.4.2 Extended Analysis
@@
Humans mostly lean towards positive decisions ($67\%$ ’yes’ answers), whereas most LLMs are more restrictive ($44\%$-$59\%$ ’yes’ answers). A manual answer analysis reveals that human annotators often decide positively despite raising several negative points. They frequently weigh the raised negative points more lightly, reserving a negative decision for extreme cases. LLMs appear to make more harsh conclusions when the accompanying textual answer contains several points of criticism. This difference in weighing weaknesses in the paper explains the comparatively low performance on decision making for the boolean nodes. Additionally, this points to the risk of a potential negativity bias of LLMs while diagnosing a paper.
@@
We turn to a language analysis of the textual answers. LLMs consistently generate longer answers ($67$ tokens on average for GPT-3.5) than humans ($55$ average tokens). Upon manual inspection, we find that the automatically generated answers to extract steps tend to split up facts exhaustively, whereas humans make broader and summarizing statements. For infer steps, LLMs often restate the inputs from prior steps, unlike human annotators. Both factors might contribute to the high factuality scores by automatic metrics because LLMs appear more specific and cover many facts, whereas humans tend to select a minimal set of facts relevant to their individual assessment process. This encourages further research on automatic evaluation metrics for this setting that can account for the reasoning context of a step.

### v2

@@ A.3.1 Interviews and Formative Study
First, there are two general types of papers in biomedicine relevant to their work: natural science papers, which investigate natural phenomena, and method papers, which propose a new approach to conducting experiments. Second, biomedical papers follow a relatively rigid surface structure with an introduction including background information, the methods section as a detailed experimental log, the results section reporting the exact measurements, and the discussion section followed by an overall conclusion. The results and discussion section can be intermingled. Third, figures and tables play a major role in the understanding and assessment of a paper. Fourth, methodological soundness and consistency of results are key indicators during text assessment. Finally, several heuristics exist to detect bad experiment design, such as the lack of controls or suspiciously high standard deviations on measurements.
@@
We conducted the interviews with each expert in isolation and oriented towards best practices of interview design by Knott et al. (2022). To structure the interview, we discussed the assessment criteria and process based on reviewing checklists of the field, starting with a general and intuitive description of the assessment process and increasing the level of detail. For the first interview, we considered the manuscript checklist by Seals and Tanaka (2000) conceptualized for physiology research. For the second interview, we use the Portland Press review checklist [^8] https://portlandpress.com/pages/peer_review_checklist (accessed 08/2024) for biomedical sciences. We opted for two different checklists to encourage different perspectives during the workflow elicitation and mitigate the impact of the interviewer’s biases.
The result of these interviews is two different step-wise manuals on how to iterate through a biomedical paper and assess its methodological soundness. We identify the there types of actions taken during assessment (read, extract, infer) which align between both manuals. While some of the subordinate steps deviate and the researchers put a different emphasize on each of these aspects, the overall assessment approach is very similar. In summary, first, the whole article is skimmed to get an understanding of its contents. Then, the experts check the contextualization with related work and background information in the introduction. Afterward, they skip to the results section, which includes figures and tables. They check the alignment of results and whether the reporting seems plausible and follows best scientific practice. Subsequently, they read the discussion in detail and analyze if the provided interpretations and conclusions align with their own readings of the results and their context in related work. Then, they verify that the method section covers all important methods and that they are described correctly. Finally, they check the conclusion section, abstract, and title to verify that they align with the findings and methodology described in the other sections. Both experts agree on the temporal ordering of these major steps during assessment.
@@
After the interviews, we consolidate the sequential manuals into one manual covering all the mentioned points by both researchers. We then interpret the temporal ordering and the subordination relationship between major and minor steps as causal dependencies to answer follow-up questions. For instance, to answer a sub-ordinate step, "Is the description of the state-of-the-art complete?" we require the input from the super-ordinate step, "How do the authors describe the state-of-the-art in the field?".
@@
First, the annotators did not finish the full workflow of $106$ steps within the three hours of the study, suggesting that the workflow should be subselected to keep the annotation practical. Second, the annotators tended to not use the causal dependencies explicitly between the steps, meaning that they answered several steps based on the paper and not referring to their own prior answers. The main obstacle for this lied in the lack of a tailored annotation interface that presents prior answers and enforces the causal structure. Third, annotators need training to get an overview of the workflow graph to avoid redundantly stating similar facts across multiple steps. An easy option to revise prior answers might also be interesting for that purpose.
@@
A.3.2 SCM for Biomedical Paper Assessment
@@
The full causal graph of the biomedical workflow is illustrated in Figure 5. Tables 4 to 6 list all workflow steps, including their type and prompt. The description and example are omitted for readability. We refer the reader to the associated codebase for the complete workflow step specifications.
@@
A.4 Data Collection
A.4.1 Underlying Paper Dataset
@@
A.4.2 Annotation Study Details
@@
Dataset Size and Generalizability
We take several measures to ensure that the findigns from our analysis (Section 4.3) and LLM experiments (Section 5) on the dataset are general and representative. As detailed beforehand, we ensure representativity of the paper sample by sampling biomedical preprints across multiple dimensions. The dataset captures roughly $4000$ reasoning steps, with three to five expert responses per paper, contributed by roughly $40$ experts in over seven hundred hours of work. We deem this amount of data sufficient for our analysis.
To investigate whether the dataset size is sufficient to make general findings on LLM performance, we perform a statistical analysis. Specifically, we investigate whether adding more papers would affect our experimental findings. To examine this empirically, we investigate a proxy: the impact of a single paper on LLM evaluation. If removing or adding one paper to the dataset does not significantly alter the results, the dataset can be considered adequate and the findings generalizable.
We test this hypothesis using statistical analysis. We focus on BERT-F1 scores as the evaluation metric. The null hypothesis states that the mean performance on the full dataset equals the mean performance on subsets excluding one paper. Only if the dataset is of sufficient size, leaving out a single paper will not induce high variance in the results. For each leave-one-paper-out subsampled dataset, we compute the mean performance and compare it to the overall mean performance computed on all papers using a single-sample t-test. To account for multiple models and therefore multiple testing, we apply Bonferroni correction. At a significance level of $0.05$ (Bonferroni-corrected to $0.0163$), the null hypothesis is not rejected; p-values exceed 0.98 across all models (specifically $0.9872$, $0.9852$, and $0.9898$). These results indicate no significant effect on performance evaluation when omitting a paper from the dataset. Thus, the dataset size is sufficient for the objectives of this study since the results are not sensitive to adding or removing a paper.
A.5 Transfer Study
To validate the generalizability of the framework, we test the transfer of the framework to case of paper assessment in natural language processing with distinct requirements. We report the detailed results here.
We employ in total $33$ PhD students specialized in NLP research to conduct a paper assessment based on the adapted workflow. This resulted in $9$ full executions (answering every item) and $24$ partial executions of the workflow. We received nearly $1000$ answers to individual reasoning steps on top of six NLP papers. To validate the workflow for the domain, we conducted an optional survey among the participants for which we received ten responses.
Rated on a scale from 1 (strongly disagree) to 7 (strongly agree), 60% of respondents assigned a score of at least 5 (above neutral) to the alignment with personal assessment practices. In free-form feedback, some of the participants noted that some steps appear to discuss overlapping paper aspects. Others noted that the provided schema worked very well for them. Overall, the transfer study results indicate that adapting the proposed approach to paper assessment is feasible with reasonable effort.
A.6 Data Insights
A.6.1 Detailed Analysis of Answer Variability
@@
A.6.2 Structural Equation Approximation from Data
@@ A.7.1 Experimental Configurations
We design two prompts for each model, one for the extract steps and one for the infer steps. Table 10 shows the prompts for LLama-3, Table 11 for Mixtral, Table 8 for GPT-3.5-turbo, and Table 9 for GPT-4o. The template parameters are substituted before inference: the parent parameter is replaced by the paper text or prior answers in the graph; the task parameter is replaced by the prompt for that step in the workflow; the description is replaced by the detailed task description of that step in the workflow; the format specifies a json object matching the expected output type (e.g. including a boolean field for boolean questions); example is substituted by the same example that annotators saw during annotation but in a suitable json format. For feedback and refinement, we use the prompts specified in Table 12 and in Table 13 for all models. Here, the true verdict is replaced by the decision and associated explanation of the task’s verdict. The graph is specified using incident encoding on integer numbered nodes; i.e. node connections are provided in the format "node x is parent of node y". The title takes the paper’s title and the answers it the list of initial answers as JSON objects including a restatement of the task prompt and dependencies for each step.
@@
Feedback Prompt
"system": "You are a specialized AI assistant focused on biomedical research quality assessment and scientific paper evaluation. You have expertise in analyzing research methodology, study design, and the logical relationships between different quality metrics.", "human": "Task: Evaluate and provide feedback on a set of dependent answers assessing the quality of the biomedical article "{title}". Context: - You are reviewing answers to a sequence of quality assessment questions - The questions form a directed acyclic graph (DAG) of dependencies - Each answer should only depend on its parent nodes in the graph - The final goal is to validate or correct the assessment to match the known true verdict *exactly* (i.e. not only the decision, but the explanation for it) Input Structure: 1. Article Title: "{title}" 2. Set of Answers: {answers} Format expected for answers: {{ "question_id": {{ "node": "question_id", "question": "question", "question_type": "type of question", "answer": json object with answer, "parent_nodes": [list of parents] }} }} 3. Dependency Graph: {graph} 4. True Final Verdict: {true_verdict} Required Analysis: 1. Consistency Check: - Verify that each answer only uses information from its parent nodes - Identify any circular reasoning or skipped dependencies 2. Logical Flow Analysis: - Trace the path from initial answers to final verdict - Identify breaks in the logical chain - Flag contradictions between connected answers 3. Verdict Alignment: - Compare the current final verdict with the true verdict - If different, identify the minimal set of answers that need revision Output Requirements: 1. Start your response with "Feedback:" 2. For each suggested revision: - Identify the specific answer(s) needing change - Explain why the change is needed - Describe how the change affects dependent answers - Provide specific suggestions for improvement 3. Prioritize changes that: - Maintain the dependency structure - Require minimal modifications - Lead to the correct final verdict Example Answer Format: Feedback: 1. Critical Issues: - [List of major inconsistencies] 2. Required Revisions: - Answer X: [Current] → [Suggested revision] Reason: [Explanation] Impact: [Effects on dependent answers] 3. Optional Improvements: - [Additional suggestions for clarity/completeness]"
Table 12: Feedback prompt used by all models during self-refinement. {…} Mark template parameters. For Mixtral the system prompt is presented as a regular human prompt. The graph is provided using incident encoding using integer indexes to describe the nodes.
Refine Prompt
"system": "You are a helpful AI assistant specializing in biomedical research and academic paper quality assessment. You are skilled at analyzing complex dependency relationships and ensuring logical consistency in multi-step evaluations.", "human": "Task: Revise a set of interdependent answers evaluating the quality of a biomedical article based on provided feedback and a known final verdict. Input Format: - title: The title of the biomedical article being evaluated: "title" - answers: JSON object containing question IDs and their current answers {answers} - graph: Directed acyclic graph in adjacency list format showing question dependencies {graph} - true_verdict: The validated final quality assessment of the article: {true_verdict} - feedback: List of specific comments on the current answers: {feedback} Dependencies: - The graph shows how questions relate: an edge from X→Y means Y’s answer should only depend on X and other parent nodes - All revised answers must maintain logical consistency with their dependent nodes - The final verdict must be *exactly* supported by the chain of revised answers; i.e. the decision and the explanation must be supported - The new final verdict should use the true verdict verbatim Constraints: 1. Maintain the original graph structure 2. Make minimal necessary changes to align with feedback 3. Ensure all answers support the true_verdict 4. Preserve the logical flow from parent to child nodes 5. Keep the original answer format unless feedback specifically suggests changes Expected Output Format: {{ "question_id": {{ "node": "question_id", "question": "question", "question_type": "type of question", "answer": json object with answer, "parent_nodes": [list of parents] }} }} Please maintain this structured format in your revised answers and ensure all changes are traceable to either the feedback or the need to align with the true verdict. Return only the JSON object containing the corrected answers with the true verdict.
Table 13: Refine prompt used by all models during self-refinement. {…} Mark template parameters. For Mixtral the system prompt is presented as a regular human prompt. The graph is provided using incident encoding using integer indexes to describe the nodes.
@@
A.7.2 Extended Results
In this section we report on more results mentioned in the main paper.
Additional Ablations and Metrics
Table 14 shows the full results including additional ablations and all metrics. Besides measuring the performance of models in the self-refinement paradigm, we test them by executing them just as LLM programs without any form of backtracking to ensure consistency to the final verdict. Furthermore, we report the F1-score both considering the individual annotators answer as ground truth and when aggregating them to a majority label.
[…]
human∗† $\textbf{0.799}_{\pm.06}$ $-0.151_{\pm.30}$ $0.151_{\pm.27}$ $-0.235_{\pm 0.34}$ 0.859 0.801
Llama3 Prg $0.752_{\pm 0.10}$ $-0.274_{\pm 0.36}$ $0.098_{\pm 0.30}$ $-0.138_{\pm 0.274}$ $0.151$ $0.170$
Mixtral Prg $0.761_{\pm 0.09}$$\textbf{-0.149}_{\pm 0.26}-$ $0.120_{\pm 0.32}$ $-0.109_{\pm 0.22}$ $0.553$ $0.559$
GPT3.5t Prg $0.759_{\pm 0.10}$ $-0.178_{\pm 0.35}$ $\textbf{0.163}_{\pm 0.37}$$\textbf{-0.102}_{\pm 0.262}-$ $0.529$ $0.531$
GPT4o Prg $0.780_{\pm 0.07}$ $-0.186_{\pm 0.30}$ $0.139_{\pm 0.35}$ $-0.138_{\pm 0.28}$ $0.718$ $0.720$
Llama3 Prgβ $0.752_{\pm.10}$ $-0.274_{\pm.36}$ $0.098_{\pm.30}$ $0.151$ $0.170$
Mixtral Prgβ $0.761_{\pm.09}$$\textbf{-0.149}_{\pm.26}-$ $0.120_{\pm.32}$ $0.553$ $0.559$
GPT3.5t Prgβ $0.750_{\pm.11}$ $-0.165_{\pm.33}$ $0.159_{\pm.37}$ $0.513$ $0.512$
GPT4o Prgβ $0.779_{\pm.07}$ $-0.196_{\pm.29}$ $0.133_{\pm.34}$ $0.696$ $0.699$
majorityio $0.854$
human∗io $0.799_{\pm.06}$ $-0.158_{\pm.29}$ $0.150_{\pm.27}$ $0.801$
Llama3io $0.786_{\pm.07}$ $-0.141_{\pm.30}$ $0.145_{\pm.35}$ $0.657$
Mixtralio $0.794_{\pm.07}$$\textbf{-0.077}_{\pm.27}-$ $0.161_{\pm.37}$ $0.822$
GPT4oio $0.795_{\pm.07}$ $-0.094_{\pm.28}$ $0.194_{\pm.40}$ 0.876
GPT4o§ $0.776_{\pm 0.07}$ $-0.154_{\pm 0.28}$ $0.188_{\pm 0.39}$ $0.843$ $0.828$
GPT4oγ $0.775_{\pm.07}$ $-0.156_{\pm.28}$ $0.187_{\pm.39}$ $0.842$ $0.827$
[…]
Error Propagation in the Graph
The stark performance gap between LLMs run as a program and human-LLM joint performance suggests error propagation as a factor. To examine, we discard the graph structure and run GPT-4o independently for each step, using the full paper as input and applying the same self-refinement loop. Results, shown in the last line of Table 3, reveal improved performance compared to GPT-4o operating as a program, though still below its performance with human oversight. This suggests that errors on earlier steps in the workflow deprive later steps of information or provide incorrect information for prediction.
A.7.3 Human Evaluation
The use of automatic evaluation in our study is a necessity, as human evaluation of all system outputs would be prohibitively expensive and prevent iterative development. While the automatic metrics used in this work are widely adopted in the community, they still remain only imperfect approximators of the real performance perceived by the users. To gain deeper insights into the relationship between our automatic metrics and human judgements for the tasks at hand, we conducted a pilot study with human evaluators.
Annotator Pool
We hire $7$ human annotators with biomedical experience through the Prolific platform [^25] https://app.prolific.com/. All annotators are paid at least $10.40$£ per hour and their payment approved regardless of their annotation performance. We require the annotators to have at least a post-graduate degree in biomedical or biochemical sciences to ensure relevant expertise. All annotators self-report being proficient in English; half of them are native speakers. All annotators indicate that they read biomedical papers regularly. We provide two abstracts of papers included in the study asking annotators to verify their expertise with respect to the specific field (synthetic biology) before taking up the task.
Setup
The objective of the human evaluation is to assess how closely the LLM-generated responses and other human answers align with the true human answer for a given workflow step. To streamline this process, we select $8$ out of the $45$ workflow steps, focusing on early-stage steps (extract steps) and critical boolean decision steps identified by their average causal effect on the final outcome (see Section 4.3). The evaluation is divided into two parts based on these step types. In the first part, annotators determine which facts extracted by the LLM are implied by the original human answer. A dedicated interface displays the predicted and the ground-truth facts. Annotators assess whether each predicted fact is implied by one, multiple, or none of the original facts. We call this task the matching task. In the second part, annotators evaluate the alignment of two explanations for boolean decisions. They assess whether the facts in the explanations contradict or align and whether similar points are raised, ignoring the decision outcome itself. We call this task the explanation alignment task.
For each workflow item, we compare the true human answer with two other human answers, the response from Mixtral (using self-refinement and paper text with human figure descriptions), and the response from GPT-4o (using self-refinement and paper text with figures as images). We choose these models because they are the top-performing LLMs. For all $8$ workflow steps, we select $3$ papers and pick the response of one annotator from the original study as the ground truth. For each item, annotators view one combination of these outputs alongside the true human answer to assess alignment. Annotation items are randomly sampled and presented in random order. Annotators first evaluate extract steps, then boolean infer steps, using provided guidelines and examples. We screen the annotator answers considering the time of annotation per item and selected items with a known correct solution.
Resulting Data
The annotation study results in $40$ annotation items of which $14$ items across the $8$ workflow steps have at least two annotations. $7$ annotators participated in the study; we filter out one annotator based on the described screening.
We measure the level of inter-annotator agreement using Krippendorff’s $\alpha$ with categorical labels for each of the tasks. For the matching task, the IAA lies at $\alpha=0.412$ computed on all pairwise comparisons of facts, for explanation alignment at $\alpha=0.444$. This moderate level of agreement is similar to other tasks aiming to identify links between two texts (e.g. implicit linking in Kuznetsov et al. (2022)) and point to the difficulty of the task similar to annotation scenarios in argumentation mining Lawrence and Reed (2020); Iskender et al. (2021). Extensive annotator training, paper context, and a rectification protocol are appropriate measures to foster consistency Klie et al. (2024). We leave this to future work because we deem the given level of consistency sufficient to estimate the relative performance differences on the task.
Evaluation Results
We aggregate the matching annotations to compute the average overlap between the facts of the predicted answer and the original human answer. We compute the Jaccard metric considering matched items as shared items for each annotator and then average over all. For explanation alignement we take the majority vote by all annotators on an item and compute the average amount of aligned explanations. Table 15 shows the results. The results closely resemble the findings based on the automatic metrics with Mixtral and GPT-4o performing nearly on-par and the human conservative baseline lagging slightly behind. GPT-4o’s performance is better on boolean decision steps than Mixtral but worse during extraction of facts.
Finally, we estimate how well the human evaluation scores align with the automatic metrics. Specifically, we compute the Pearson correlation to determine whether there is a linear relationship with the automatic metrics. Table 16 shows the correlation coefficients for the two evaluation scenarios. BERT-F1 overall correlates moderately with human judgement in both scenarios. As a general purpose text similarity metric, BERT-F1 captures the answer similarity well regardless of the underlying workflow step. SummaC, on the other hand, appears inappropriate for matching evaluation (weak negative correlation), but well-suited for alignment. The TRUE score show overall low to moderate alignment with human judgement. In conclusion, the used automatic metrics capture human judgement to some extend and can serve as a solid estimate for performance. The differences in correlation to human judgement for the two different task types in the workflow suggest that a more fine-grained, step-specific evaluation is a promising direction of future research.
matching alignment avg
human $0.339$ $0.5$ $0.42$
Mixtral $0.477$ $0.667$ $0.572$
GPT-4o $0.351$ $0.75$ $0.551$
Table 15: Based on the human evaluation data we compute the Jaccard metric on facts for matching items and compute the average amount of aligned explanations. To aggregate into a single metric, we take the average of both.
matching alignment
BERT-F1 $0.37$ $0.54$
SummaC $-0.22$ $0.55$
TRUE $0.30$ $0.20$
Table 16: Spearman’s $\rho$ comparing the human evaluation scores with the automatic metrics for the two scenarios.
A.7.4 Extended Analysis
@@
Humans mostly lean towards positive decisions ($67\%$ ’yes’ answers), whereas most LLMs are more restrictive ($44\%$-$59\%$ ’yes’ answers). A manual answer analysis reveals that human annotators often decide positively despite raising several negative points. They frequently weigh the raised negative points more lightly, reserving a negative decision for extreme cases. LLMs appear to make more harsh conclusions when the accompanying textual answer contains several points of criticism. This difference in weighing weaknesses in the paper explains the comparatively low performance on decision making for the boolean nodes. Additionally, this points to the risk of a potential negativity bias of LLMs while assessing a paper.
@@
We turn to a language analysis of the textual answers. LLMs consistently generate longer answers ($67$ tokens on average for GPT-3.5) than humans ($55$ average tokens). Upon manual inspection, we find that the automatically generated answers to extract steps tend to split up facts exhaustively, whereas humans make broader and summarizing statements. For infer steps, LLMs often restate the inputs from prior steps, unlike human annotators. Both factors might contribute to the high factual alignment by automatic metrics because LLMs appear more specific and cover many facts, whereas humans tend to select a minimal set of facts relevant to their individual assessment process. This encourages further research on automatic evaluation metrics for this setting that can account for the reasoning context of a step.